The average ACT score follows a normal distribution with a mean of 21.1 and a standard deviation of 5.1 what is the probability that the mean IQ score of 50 randomly selected people will be more than 23 ?
Answer #1

The sampling distribution of 50 randomly selected people will follow a normal distribution with:

Mean = 21.1

s = 5.1/√50 = 0.72

We have to find: P(X > 23) = P( z > (23 - 21.1)/0.72 ) = P(z > 2.64)

From the z-table,

P(z > 2.64) = 1 - P(z < 2.64) = 1 - 0.9959 = 0.0041